I know the answer to this equation. It is 6, but I don't know how to show my work. I don't know how to show how I got to that answer. 17+[(8w-3)/9]=22

17+[8w-3/9]=22
17+-1/3=22
132/6
6=6

Found 2 solutions by askmemath, jacobskayla_lee_633:
Answer by askmemath(368)

Multiplying by 9 throughout
153 +8w - 3 = 198
150+8w = 198
8w = 198-150
8w = 48
w = 6

Answer by jacobskayla_lee_633(5)